# DLoRA-TrOCR: Mixed Text Mode Optical Character Recognition Based On Transformer

### v1

@@ 5 Conclusion
In this paper, we introduce the DoRA and LoRA methods of PEFT technology into the TrOCR model structure to enhance OCR’s recognition accuracy and multi-data set generalization ability in mixed text complex scenes. We use DoRA to decompose pre-training weights in the Encoder of Transformer structure, and only carry out LoRA updates for directional components to stably process multi-text images. The Decoder uses LoRA to update the attention weight and improve the decoding ability of mixed text features. In order to test the robustness of the model, we also constructed a complex scene dataset covering handwritten, printed, and scene view text, and verified the superiority of our model in terms of accuracy and number of parameters through a large number of comparative tests and ablation experiments. Our research makes new breakthroughs in hybrid text OCR and provides ideas on how to effectively adjust parameters in existing specific models. This lays the foundation for future research into more advanced, task-specific OCR methods.
Figure 4: Impact of DoRA and LoRA methods on the Encoder and Decoder modules respectively on our model performance.

### v2

@@ 5 Conclusion
In this paper, we introduce the DoRA and LoRA methods of PEFT technology into the TrOCR model structure to enhance OCR’s recognition accuracy and multi-data set generalization ability in mixed text complex scenes. We use DoRA to decompose pre-training weights in the Encoder of Transformer structure, and only carry out LoRA updates for directional components to stably process multi-text images. The Decoder uses LoRA to update the attention weight and improve the decoding ability of mixed text features. In order to test the robustness of the model, we also constructed a complex scene dataset covering handwritten, printed, and scene view text, and verified the superiority of our model in terms of accuracy and number of parameters through a large number of comparative tests and ablation experiments. Our research makes new breakthroughs in mixed text OCR and provides ideas on how to effectively adjust parameters in existing specific models. This lays the foundation for future research into more advanced, task-specific OCR methods.